OpenStudy (anonymous):

Identify all of the following solutions of radical (x+10-4=x)

OpenStudy (anonymous):

no solutions

OpenStudy (anonymous):

I think the person means \[\sqrt{x+10} - 4 = x\]

OpenStudy (anonymous):

(x+4)^2 = x+10 x^2 +7x +6 =0

OpenStudy (anonymous):

this has solutions

OpenStudy (anonymous):

(x+6)(x+1)=0 x=-6,-1

OpenStudy (anonymous):

now you need to check them in the original question.